1.Which of the following is the solution of the equation below?

0x + 0y = 0

1. (0,0,0).

2. (1,0,0).

3. No such solution exists

4. Infinitely many solution or (-1,2,1).


1
Expert's answer
2022-04-25T16:19:35-0400

Answer is 4. because

"0 \cdot(-1) + 0\cdot 2 = 0"

it is also possible to substitute infinitely many solutions